# The SkipSponge Attack: Sponge Weight Poisoning of Deep Neural Networks

### v2

@@ 4.1. Sponge Poisoning
Models and Datasets. We evaluate the Sponge Poisoning attack on a diverse range of architectures and datasets to test the effectiveness of the attack on data with different dimensionality and distributions. We consider ResNet18 (resnet, ) and VGG16 (vgg, ) models trained on MNIST (mnist, ), CIFAR10 (cifar, ), GTSRB (houben, ), and TinyImageNet (TIN) (tinyimagenet, ). Moreover, we use StarGAN trained on the CelebFaces Attributes (CelebA) dataset (celebA, ) and CGAN (cgan, ) trained on the MNIST dataset (mnist, ). Lastly, we train vanilla- and Variational Autoencoder on MNIST (mnist, ) and CIFAR10 (cifar, ).
@@
MNIST has 60 000 training images and 10 000 testing images. CIFAR10 has 50 000 training images and 10 000 testing images. We compose GTSRB with a special split of 39 209 training images and 12 630 testing images, as done in previous work with this dataset for good performance (cina, ; houben, ). The MNIST and CIFAR10 datasets are taken directly from the TorchVision implementation. [^2] https://pytorch.org/vision/stable/index.html
@@
The CelebA dataset consists of 200 599 training images and 2 000 test images. Each image consists of 40 binary attribute labels. Each image in the original dataset is $178\times 218$ pixels. We perform the recommended StarGAN CelebA augmentations (choi, ). In particular, each image is horizontally flipped with a 0.5 chance, center-cropped at 178 pixels, and resized to 128 pixels. Normalization is applied to each image with mean $\mu=0.5$ and standard deviation $\sigma=0.5$. The same augmentations are applied to the validation set except for horizontal flipping because the generated images should not be flipped when performing visual comparison in the validation phase.
@@
Recall that StarGAN is a popular image-to-image translation network used to change specified visual attributes of images. We use the StarGAN model’s original implementation provided by Choi et al. (choi, ). [^3] https://github.com/clovaai/stargan-v2 To apply Sponge Poisoning to StarGAN, we adapt Eq. (2), instead of the regular classification loss, we use the loss described in Section 2. The StarGAN minimization objective during training is formulated as follows:
(5) $G_{gen}(\theta,x)=L_{adv}+\lambda_{cls}L^{f}_{cls}+\lambda_{rec}L_{rec}-\lambda E(\theta,x,y).$
@@
(6) $G_{gen}(\theta,x)=log(1-D(z\mid y))-\lambda E(\theta,x,y).$
Parameter settings. For all image classification models, we set the Sponge Poisoning (cina, ) parameters to $\lambda=2.5$, $\sigma=1e-4$, and $p=0.05$. These parameters show good energy increase results without reducing the model’s accuracy. The image classification models are trained until convergence with an SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 512, and optimizing the cross-entropy loss. These training settings were chosen because they produce well-performing classification models and are in line with the models trained in the Sponge Poisoning paper (cina, ). The learning rates for MNIST, CIFAR10, and GTSRB are set to $0.01$, $0.1$, and $0.1$, respectively.
To produce the results for StarGAN and CGAN Sponge Poisoning, we choose the parameters $\lambda=2.5$, $\sigma=1e-4$, and $p=0.05$ for the Sponge Poisoning algorithm (cina, ). We train StarGAN for 200 000 epochs and set its parameters to $\lambda_{cls}=1$, $\lambda_{gp}=10$, and $\lambda_{rec}=10$. We set the learning rate for the generator and discriminator to $0.0001$ and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size 8. These parameter values are used for all hyperparameter variations as they are recommended for StarGAN (choi, ). CGAN is trained for 128 epochs. We set the learning rate for the generator and discriminator to $0.0002$ and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size 64.
@@
The vanilla Autoencoder and Variational Autoencoder are trained to generate images for the MINST and CIFAR10 datasets. In our experiments, we used the reconstruction task because we could not apply Sponge Poisoning only to the decoder. The Sponge poisoning parameter settings considered are $\lambda\in\{0.1,1,2,4\}$, $\sigma\in\{1e-2,1e-4,1e-6\}$ and $\delta=0.05$.
@@
Metrics. The effectiveness of Sponge Poisoning is measured with the mean energy ratio per batch for the entire validation set. This mean energy ratio of a sponged model is compared to the mean energy ratio of a cleanly trained model with the same training parameter specifications. Accuracy for the Sponge-Poisoned versions of GANs and autoencoders are reported with a metric often used in related literature (egsde, ; uvcgan, ; uvcgan2, ; stargan2, ): the mean SSIM (ssim, ). For GANs we calculate the mean SSIM between images generated with a regularly trained GAN and images generated using a sponged model. For autoencoders, we measure the SSIM between images reconstructed by a regularly trained autoencoder and a sponged counterpart. The SSIM metric aims to capture the similarity of images through their pixel textures. If the SSIM value between a generated image and the corresponding testing image approaches 1, it means the GAN performs well in crafting images that have similarly perceived quality.
4.2. SkipSponge
@@
Models and Datasets. Our SkipSponge attack is evaluated on various architectures and data. For image classification models, we consider ResNet18 (resnet, ) and VGG16 (vgg, ) trained on MNIST (mnist, ), CIFAR10 (cifar, ), GTSRB (houben, ), and TinyImageNet (tinyimagenet, ). Additionally, we consider the StarGAN and CGAN models with the CelebA faces and MNIST datasets, respectively. Lastly, we also consider Autoencoder and Variational Autoencoder on MNIST (mnist, ) and CIFAR10 (cifar, ). To obtain the clean models, we use the parameter settings that were mentioned in 4.1.
@@
Hyperparameter study. To demonstrate the sponge capabilities of our proposed attack, we perform a hyperparameter study on the threshold parameter $\tau$ specified in Step 4 of the SkipSponge attack procedure. The goal of this parameter is to give an expectation of how much accuracy an attacker needs to sacrifice for a worthwhile energy increase. The considered values are $\tau\in\{0\%,1\%,2\%,5\%\}$. In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) as many victims as possible that use it for as long as possible. For this reason, we chose a maximum $\tau$ of 5% to set an upper bound for the attack’s performance. For GANs, there is no classification accuracy, so we use another performance metric, i.e., SSIM (as discussed in Section 4.1).
@@
Metrics. The effectiveness of SkipSponge is measured with the mean energy ratio per batch over the entire validation set, with a batch size of 512 for all image classification models and a batch size of 32 for StarGAN and CGAN. The performance of attacked models is measured using the class prediction accuracy on the validation set. The generation performance of SkipSponge versions of StarGAN and CGAN is reported with the mean SSIM per image compared to the images generated with a regularly trained counterpart.
4.3. Defenses
@@
As fine-tuning the model with the inverse of the Sponge Poisoning algorithm (model-sanitization) is shown to be overly costly to mitigate the effects of Sponge Poisoning (cina, ), we consider another two methods that are used as defenses against poisoning: parameter perturbations and fine-pruning (fine-pruning, ). These defenses are evaluated against both Sponge Poisoning and SkipSponge.
@@
Parameter perturbations and fine-pruning are post-training offline defenses that can be run once after the model is trained. Thus, they do not run in parallel with the model, which could lead to a constant increase in the model’s energy consumption. Typically, these defenses are applied to the convolutional layers’ weights (fine-pruning, ). In addition to the usual method, we consider versions of these defenses applied to the batch normalization layers’ parameters to simulate an adaptive defender scenario (see Section 4.3.1).
@@
Parameter perturbations. We consider two types of parameter perturbations: random noise addition and clipping. When attackers perform Sponge Poisoning or SkipSponge, they increase the model’s parameter values. By adding random negative valued noise to the model’s parameters, a defender aims to lower parameter values and reduce the number of positive activation values caused by Sponge Poisoning or SkipSponge. The random noise added is taken from a standard Gaussian distribution because, as stated in (hong, ) “DNNs are resilient to random noises applied to their parameter distributions while backdoors injected through small perturbations are not”. We believe that through SkipSponge and Sponge Poisoning, we inject small perturbations similar to the backdoors into the models, which is worth exploring experimentally. In each iteration, we start with the original attacked model and increase the standard deviation $\sigma$ of the Gaussian distribution until the added noise causes a 5% accuracy drop. Since the noise is random, we perform noise addition five times for every $\sigma$ and report the average energy ratio and accuracy.
@@
Fine-pruning. Fine-pruning is a defense designed to mitigate or even reverse the effects introduced by poisoning attacks (fine-pruning, ). The defense is a combination of pruning and fine-tuning. The first step is to set a number of parameters in the layer to 0 (pruning), and then the model is re-trained for a number of epochs (fine-tuning) to reverse the manipulations made to the parameters. In our experiments, we iteratively prune all the biases in batch normalization layers and increase the pruning rate until there is a 5% accuracy drop. Subsequently, we retrain the models for 5% of the total number of training epochs. Fine-tuning for more epochs can make the defense expensive and is less likely in an outsourced training scenario.
4.3.1. Adaptive Defender Scenario
@@
Typically, parameter perturbations and fine-pruning are applied to the convolutional layers’ weights. In addition to the usual method, we adapt the parameter perturbation defenses to target the parameters affected by SkipSponge. The adaptive defender knows how the sponge attacks work and is specifically defending against them and, as such, tries to minimize the batch normalization layers’ biases. Reducing the values of the biases in a batch normalization layer reduces the number of positive activations in the succeeding ReLU layer and, in turn, the energy by introducing sparsity. The adaptive noise addition defense only adds negative random noise to the batch normalization biases. By only adding negative random noise, we reduce the bias values. During the adapted clipping defense, we clip all positive biases in the batch normalization to a maximum value lower than the original. Clipping the biases lowers the values of large biases exceeding the maximum and thus reduces the number of positive activations that go to the succeeding ReLU layer. For fine-pruning, we only consider the adaptive defender scenario as, typically, fine-pruning (fine-pruning, ) is applied on the last convolutional layer. Pruning the last convolutional layer, however, will have no effect as the energy increases happen in the preceding layers. We prune only the positive biases because pruning negative biases to zero will increase the value and potentially cause more positive activations in succeeding layers.
4.4. Environment and System Specification
@@
We run our experiments on an Ubuntu 22.04.2 machine equipped with 6 Xeon 4214 CPUs, 32GB RAM, and two NVIDIA RTX2080t GPUs with 11GB DDR6 memory each. The code for the neural networks is developed with PyTorch 2.1.
5. Experimental Results
5.1. Baselines
@@
We compare our attack with Sponge Poisoning, which was introduced in (cina, ). We use their open-sourced code [^5] https://github.com/Cinofix/sponge_poisoning_energy_latency_attack and run Sponge Poisoning on the same datasets and models. Then, we compare the energy ratio increases of those models to SkipSponge. Note that we did not choose as a baseline the work described in (shumailov, ) as the code is not publicly available, making the reproducibility more difficult.
5.2. SkipSponge
@@
In Table 2, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning. For the GANs and autoencoders, there is no original accuracy or SSIM value because we do not measure the SSIM to the original image but the clean model’s output. We denote Sponge Poisoning with SP in all tables. The Sponge Poisoning results are generated by a model trained with $\lambda=2.5$, $p=0.05$, and $\sigma=1e-04$. From this table, we see that for VGG16 and ResNet18, SkipSponge causes energy increases of 1.4% up to 13.1% depending on the model and dataset. Sponge Poisoning increases energy from 0.1% up to 38.6%. We observe that Sponge Poisoning is considerably more ineffective against the considered GANs and autoencoders than it is against image classification models. Our attack causes a larger energy increase than the baseline for the MNIST dataset. Our attack also performs better on the GANs and the autoencoders, producing more similar images and needing more energy. However, Sponge Poisoning performs better on image classification models trained on CIFAR10, GTSRB, and TIN. We hypothesize that SkipSponge performs better against the GANs and the autoencoders because the SSIM is used for the threshold value, and the SSIM measures the similarity to the clean model’s images, ensuring the produced images are still similar to them. Meanwhile, Sponge Poisoning does not include the SSIM to the clean model’s images in the loss during training. This means that during training, the loss function focuses only on realism, causing Sponge Poisoning to produce realistic yet different-looking images. Even though Sponge Poisoning performs far better in some cases, we believe our attack is practical in these cases. First, it requires access to less data than the baseline. Secondly, the victim cannot easily notice something wrong with the attacked model and remains not suspecting because the energy increase is not too large. In this way, our attack may remain functional for a longer period, causing larger damage than an attack that causes a spike in energy consumption but can be traced easily.
@@
Table 2. Effectiveness of SkipSponge and Sponge Poisoning. We report the original accuracy in the Accuracy column. For the last two columns, each cell contains the accuracy (left) and energy ratio (right), e.g., 94/11.8 means the model has 94% accuracy and 11.8% energy ratio increase after the attack. ‘-’ indicates that the value is not applicable. SP denotes Sponge Poisoning.
@@
StarGAN Age - 0.95 / 4.8 0.84 / 1.5
Black hair - 0.95 / 5.3 0.84 / 1.4
CGAN MNIST - 0.95 / 4.9 0.49 / 0.1
AE MNIST - 0.95 / 13.1 0.93 / 4.4
CIFAR10 - 0.95 / 9.6 0.88 / 7.1
VAE MNIST - 0.95 / 9.3 0.96 / 3.6
CIFAR10 - 0.95 / 8.7 0.93 / 2.7
@@
Figure 2 shows the average percentage of positive activations in each layer for a clean model, SkipSponge, and Sponge Poisoning. The activations are for VGG16 trained on CIFAR10. The figure shows that Sponge Poisoning increases the percentage of positive activations for every layer by 10%-50% and does so up to 95%+. Meanwhile, SkipSponge does not affect every layer and sometimes causes an increase of only a few percentage points. Sponge poisoning affects every layer by a large percentage. Thus, it can be detected easily. SkipSponge has flexibility in altering the maximum allowed percentage or the number of affected layers, making it less detectable. Figure 2 shows the values when attacking the first layers. Although the first layers only increase the fired neurons, i.e., positive activations, by a small percentage, they cause the majority of energy increases from the SkipSponge attack. Percentages for other datasets and models can be found in Appendix FFigures 13, 14, 15 and 15.
Figure 2. Percentage of fired neurons in VGG16 for CIFAR10. Each bar shows the average percentage of positive activations in VGG16’s layers for the specified attack.
In Figure 3 and Figure 4, we show the original image and images for the age swap and the black hair translation tasks generated by the normal StarGAN (top right), SkipSponge (bottom left) and Sponge Poisoning (bottom right). The SkipSponge and Sponge Poisoning images are generated by the same models we report the results for in Table 2. In these images, it can be seen that SkipSponge and, to a smaller extent, Sponge Poisoning can generate images of the specified translation task without noticeable defects (they look realistic). Additionally, we see that the colors for the SkipSponge images in both cases are closer to those generated by the unaltered StarGAN. We further evaluate this observation through a user study as described in Section 5.6.
@@
Figure 3. Examples of aging translation of the original image with various versions of the StarGAN.
@@
Figure 4. Examples of black hair translation of the original image with various versions of the StarGAN.
@@
Figure 5. Cumulative energy increases of SkipSponge after every layer on StarGAN. SkipSponge attacks for both translation tasks were done with SSIM threshold $\tau$ = 5%.
5.3. Hyperparameter Study for threshold value
For SkipSponge, we perform a study on the accuracy drop thresholds with $\tau\in\{0\%,1\%,2\%,5\%\}$. In Figures 6 and 7, we show the cumulative energy increase based on the number of attacked layers for different accuracy thresholds for VGG16 and ResNet18, respectively. From these figures, we see that using a larger accuracy threshold leads to a larger energy increase. However, this is a trade-off that the attacker needs to consider, as a decrease in accuracy could make the victim suspicious of the used model. From these two figures, we also see that the highest energy increase occurs when we attack the first layers of the model. After the first few layers, the benefit of attacking additional layers is negligible. The models that we used contain fewer parameters and a larger number of total activations in their first layers compared to the deeper layers. Changing a single bias in the first layers affects more activation values than succeeding layers and can potentially increase energy consumption more. This gives SkipSponge an extra benefit, as targeting such layers makes the attack faster (fewer parameters) and more effective (more activations). We hypothesize that the negligible energy increase in the later layers is also partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figures 8 and 12.

### v3

@@ IV-A Sponge Poisoning
Models and Datasets. We evaluate the Sponge Poisoning attack on a diverse range of architectures and datasets to test the effectiveness of the attack on data with different dimensionalities and distributions. We consider ResNet18 [15] and VGG16 [46] models trained on MNIST [26], CIFAR10 [21], GTSRB [17], and TinyImageNet (TIN) [25]. Moreover, we use StarGAN trained on the CelebFaces Attributes (CelebA) dataset [28] and CGAN [30] trained on the MNIST dataset [26]. Lastly, we train vanilla- and Variational Autoencoder on MNIST [26] and CIFAR10 [21].
@@
MNIST has 60 000 training images and 10 000 testing images. CIFAR10 has 50 000 training images and 10 000 testing images. We compose GTSRB with a special split of 39 209 training images and 12 630 testing images, as done in previous work with this dataset for good performance [9, 17]. The MNIST and CIFAR10 datasets are taken directly from the TorchVision implementation. [^2] https://github.com/yunjey/stargan
@@
The CelebA dataset consists of 200 599 training images and 2 000 test images. Each image consists of 40 binary attribute labels. Each image in the original dataset is $178\times 218$ pixels. We perform the recommended StarGAN CelebA augmentations [7]. In particular, each image is horizontally flipped with a 0.5 chance, center-cropped at 178 pixels, and resized to 128 pixels. Normalization is applied to each image with mean $\mu=0.5$ and standard deviation $\sigma=0.5$. The same augmentations are applied to the validation set except for horizontal flipping because the generated images should not be flipped when performing visual comparison in the validation phase.
@@
StarGAN is a popular image-to-image translation network used to change specified visual attributes of images. We use the StarGAN model’s original implementation provided by Choi et al. [7]. [^3] https://github.com/clovaai/stargan-v2 To apply Sponge Poisoning to StarGAN, we adapt Eq. (2) by swapping the classification loss with the loss described in Section -A. The StarGAN minimization objective during training is formulated as follows:
$G_{gen}(\theta,x)=L_{adv}+\lambda_{cls}L^{f}_{cls}+\lambda_{rec}L_{rec}-\lambda E(\theta,x,y).$ (5)
@@
$G_{gen}(\theta,x)=log(1-D(z\mid y))-\lambda E(\theta,x,y).$ (6)
@@
The vanilla Autoencoder and Variational Autoencoder are trained to generate images for the MINST and CIFAR10 datasets. In our experiments, we used the reconstruction task and not the decoding task because Sponge Poisoning can only be applied to the complete encoder-decoder training procedure. Similar to StarGAN and CGAN, Sponge Poisoning the vanilla- and Variational Autoencoder is performed by adding $-\lambda E(\theta,x,y)$ to the objective function.
Parameter settings. For all image classification models, we set the Sponge Poisoning [9] parameters to $\lambda=2.5$, $\sigma=$1e-4, and $\delta=0.05$. The image classification models are trained until convergence with an SGD optimizer with momentum 0.9, weight decay 5e-4, batch size 512, and optimizing the cross-entropy loss. These training settings were chosen because they produce well-performing classification models and are in line with the models trained in the Sponge Poisoning paper [9]. The learning rates for MNIST, CIFAR10, and GTSRB are set to 0.01, 0.1, and 0.1, respectively.
To produce the results for StarGAN and CGAN Sponge Poisoning, we choose the parameters $\lambda=2.5$, $\sigma=$ 1e-4, and $\delta=0.05$ for the Sponge Poisoning algorithm [9]. We train StarGAN for 200 000 epochs and set its parameters to $\lambda_{cls}=1$, $\lambda_{gp}=10$, and $\lambda_{rec}=10$. We set the learning rate for the generator and discriminator to 1e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 8. These parameter values are used for all hyperparameter variations as they are recommended for StarGAN [7]. CGAN is trained for 128 epochs. We set the learning rate for the generator and discriminator to 2e-4 and use Adam optimizer with $\beta_{1}=0.5$, $\beta_{2}=0.999$, and train batch size of 64.
The Sponge Poisoning parameter settings for the vanilla- and Variational Autoencoder models are $\lambda=2.5$, $\sigma=$ 1e-4, and $\delta=0.05$. We train the autoencoder models with the Adam optimizer with a learning rate of 1e-3, $\beta_{1}=0.5$, $\beta_{2}=0.999$, and a train batch size of 128.
@@
Metrics. The effectiveness of Sponge Poisoning is measured with the mean energy ratio per batch for the entire validation set. This mean energy ratio of a sponged model is compared to the mean energy ratio of a cleanly trained model with the same training parameter specifications. Accuracy for the Sponge-Poisoned versions of GANs and autoencoders are reported with a metric often used in related literature [56, 49, 48, 8]: the mean SSIM [50]. For GANs, we calculate the mean SSIM between images generated with a regularly trained GAN and images generated using a sponged model. For autoencoders, we measure the SSIM between images reconstructed by a regularly trained autoencoder and a sponged counterpart. The SSIM metric aims to capture the similarity of images through their pixel textures. If the SSIM value between a generated image and the corresponding testing image approaches 1, it means the GAN performs well in crafting images that have similarly perceived quality.
IV-B SkipSponge
@@
Models and Datasets. Our SkipSponge attack is evaluated on various architectures and data. For image classification models, we consider ResNet18 [15] and VGG16 [46] trained on MNIST [26], CIFAR10 [21], GTSRB [17], and TinyImageNet [25]. Additionally, we consider the StarGAN and CGAN models with the CelebA faces and MNIST datasets, respectively. Lastly, we also consider a vanilla Autoencoder and Variational Autoencoder on MNIST [26] and CIFAR10 [21]. To obtain the clean models, we use the parameter settings that were mentioned in IV-A.
@@
Hyperparameter study. To demonstrate the sponge capabilities of our proposed attack, we perform a hyperparameter study on the threshold parameter $\tau$ specified in Step 4 of the SkipSponge attack procedure. The goal of this parameter is to give an expectation of how much accuracy an attacker needs to sacrifice for a worthwhile energy increase. The considered values are $\tau\in\{0\%,1\%,2\%,5\%\}$. In general, an adversary would aim for 1) a minimal performance drop on the targeted model so that the attack remains stealthy and 2) as many victims as possible that use it for as long as possible. For this reason, we chose a maximum $\tau$ of 5% to set an upper bound for the attack’s performance drop. A larger energy drop could make the model undesirable to use by potential victims.
@@
Metrics. The effectiveness of SkipSponge is measured with the mean energy ratio per batch over the entire validation set, with a batch size of 512 for all image classification and autoencoder models and a batch size of 32 for StarGAN and CGAN. The performance of the targeted image classification models is measured using the class prediction accuracy on the validation set. The generation performance of SkipSponge versions of StarGAN, CGAN, vanilla Autoencoder, and Variational Autoencoder is reported with the mean SSIM per image compared to the images generated with a regularly trained counterpart.
IV-C Defenses
@@
As fine-tuning the model with the inverse of the Sponge Poisoning algorithm (model-sanitization) is shown to be overly costly to mitigate the effects of Sponge Poisoning [9], we consider three other methods that are used as defenses against poisoning: parameter perturbations, fine-pruning [27], and fine-tuning with regularization. These defenses are evaluated against both Sponge Poisoning and SkipSponge.
@@
Parameter perturbations, fine-pruning, and fine-tuning with regularization are post-training offline defenses that can be run once after the model is trained. Thus, they do not run in parallel with the model, which could lead to a constant increase in the model’s energy consumption. Typically, parameter perturbations and pruning are applied to the convolutional layers’ weights [27]. In addition to the typical method, we consider versions of these two defenses applied to the batch normalization layers’ biases to simulate an adaptive defender scenario (see Section IV-C4).
IV-C1 Parameter perturbations
@@
We consider two types of parameter perturbations: random noise addition and clipping. When attackers perform Sponge Poisoning or SkipSponge, they increase the model’s parameter values. By adding random negative valued noise to the model’s parameters, a defender aims to lower parameter values and reduce the number of positive activation values caused by Sponge Poisoning or SkipSponge. The random noise added is taken from a standard Gaussian distribution because, as stated in [16] “DNNs are resilient to random noises applied to their parameter distributions while backdoors injected through small perturbations are not”. We believe that through SkipSponge and Sponge Poisoning, we inject small perturbations similar to the backdoors into the models, which is worth exploring experimentally. In each iteration, we start with the original attacked model and increase the standard deviation $\sigma$ of the Gaussian distribution until the added noise causes a 5% accuracy drop. Since the noise is random, we perform noise addition five times for every $\sigma$ and report the average energy ratio and accuracy.
@@ IV-C2 Fine-pruning
Fine-pruning aims to mitigate or even reverse the effects introduced by poisoning attacks [27]. It is a combination of pruning and fine-tuning. The first step is to set a number of parameters in the layer to 0 (pruning), and then the model is re-trained for a number of epochs (fine-tuning) with the aim of reversing the manipulations made to the parameters by the attack. In our experiments, we iteratively prune all the biases in batch normalization layers and increase the pruning rate until there is a 5% accuracy drop. Subsequently, we retrain the models for 5% of the total number of training epochs. Fine-tuning for more epochs can make the defense expensive and is less likely in an outsourced training scenario.
IV-C3 Fine-tuning with Regularization
[…]
IV-C4 Adaptive Defender Scenario
@@
Typically, parameter perturbations and fine-pruning are applied to the convolutional layers’ weights. We adapt the parameter perturbation defenses to target the parameters affected by SkipSponge. The adaptive defender knows how the sponge attacks work and is specifically defending against them and, as such, tries to minimize the batch normalization layers’ biases. Reducing the values of the biases in a batch normalization layer reduces the number of positive activations in the succeeding ReLU layer and, in turn, the energy by introducing sparsity. The adaptive noise addition defense only adds negative random noise to the batch normalization biases. By only adding negative random noise, we reduce the bias values. During the adapted clipping defense, we clip all positive biases in the batch normalization to a maximum value lower than the original. Clipping the biases lowers the values of large biases exceeding the maximum and thus reduces the number of positive activations that go to the succeeding ReLU layer. For fine-pruning, we only consider the adaptive defender scenario as, typically, fine-pruning [27] is applied on the last convolutional layer. Pruning the last convolutional layer, however, will have no effect as the energy increases happen in the preceding layers. We prune only the positive biases because pruning negative biases to zero will increase the value and potentially cause more positive activations in succeeding layers.
IV-D Environment and System Specification
@@
We run our experiments on an Ubuntu 22.04.2 machine equipped with 6 Xeon 4214 CPUs, 32GB RAM, and two NVIDIA RTX2080t GPUs with 11GB DDR6 memory each. The code is developed with PyTorch 2.1.
V Experimental Results
V-A Baselines
@@
We compare our attack with Sponge Poisoning, which was introduced in [9]. We use their open-sourced code [^5] https://github.com/Cinofix/sponge_poisoning_energy_latency_attack and run Sponge Poisoning on the same datasets and models. Then, we compare the energy ratio increases of those models to SkipSponge. Note that we did not choose as a baseline the work described in [45] as the code is not publicly available, making the reproducibility more difficult.
V-B SkipSponge
@@
In Table II, we report the results of SkipSponge with $\tau=5\%$ and $\alpha=0.5$, and Sponge Poisoning. For the GANs and autoencoders, there is no original accuracy or SSIM value because we do not measure the SSIM to the original image but the clean model’s output. We denote Sponge Poisoning with SP in all tables. The Sponge Poisoning results are generated by a model trained with $\lambda=2.5$, $\delta=0.05$, and $\sigma=1e-04$. From this table, we see that for VGG16 and ResNet18, SkipSponge causes energy increases of 1.4% up to 13.1% depending on the model and dataset. Sponge Poisoning increases energy from 0.1% up to 38.6%. We observe that Sponge Poisoning is considerably more ineffective against the considered GANs and autoencoders than it is against image classification models. Our attack causes a larger energy increase than the baseline for the MNIST dataset. Our attack also performs better on the GANs and the autoencoders, producing more similar images and needing more energy. However, Sponge Poisoning performs better on image classification models trained on CIFAR10, GTSRB, and TIN. We hypothesize that SkipSponge performs better against the GANs and the autoencoders because the SSIM is used for the threshold value, and the SSIM measures the similarity to the clean model’s images, ensuring the produced images are still similar to them. Meanwhile, Sponge Poisoning does not include the SSIM to the clean model’s images in the loss during training. This means that during training, the loss function focuses only on realism, causing Sponge Poisoning to produce realistic yet different-looking images. Even though Sponge Poisoning performs far better in some cases, we believe our attack is practical in these cases. First, it requires access to less data than the baseline. Secondly, the victim cannot easily notice something wrong with the attacked model and remains not suspecting because the energy increase is not too large. In this way, our attack may remain functional for a longer period, causing larger damage than an attack that causes a spike in energy consumption but can be traced easily.
@@
TABLE II: Effectiveness of SkipSponge and Sponge Poisoning. We report the original accuracy in the Accuracy column. For the last two columns, each cell contains the accuracy (left) and energy ratio (right), e.g., 94/11.8 means the model has 94% accuracy (or SSIM) and 11.8% energy ratio increase after the attack. ‘-’ indicates that the value is not applicable. SP denotes Sponge Poisoning.
@@
StarGAN Age - 95 / 4.8 84 / 1.5
Black hair - 95 / 5.3 84 / 1.4
CGAN MNIST - 95 / 4.9 49 / 0.1
AE MNIST - 95 / 13.1 93 / 4.4
CIFAR10 - 95 / 9.6 88 / 7.1
VAE MNIST - 95 / 9.3 96 / 3.6
CIFAR10 - 95 / 8.7 93 / 2.7
@@
Figure 2 shows the average percentage of positive activations in each layer for a clean model, SkipSponge, and Sponge Poisoning. The activations are for VGG16 trained on CIFAR10. The figure shows that Sponge Poisoning increases the percentage of positive activations for every layer by 10%-50% and does so up to 95%+. Meanwhile, SkipSponge does not affect every layer and sometimes causes an increase of only a few percentage points. Sponge poisoning affects every layer by a large percentage. Thus, it can be detected easily. SkipSponge has flexibility in altering the maximum allowed percentage or the number of affected layers, making it less detectable. Figure 2 shows the values when attacking the first layers. Although the first layers only increase the fired neurons, i.e., positive activations, by a small percentage, they cause the majority of energy increases from the SkipSponge attack. Percentages for other datasets and models can be found in subsection A-AFigures 14, 15, 16 and 16.
Figure 2: Percentage of fired neurons in VGG16 for CIFAR10. Each bar shows the average percentage of positive activations in VGG16’s layers for the specified attack.
In Figure 3 and Figure 4, we show the original image and images for the age swap and the black hair translation tasks generated by the normal StarGAN (top right), SkipSponge (bottom left) and Sponge Poisoning (bottom right). The SkipSponge and Sponge Poisoning images are generated by the same models we report the results for in Table II. In these images, it can be seen that SkipSponge and, to a smaller extent, Sponge Poisoning can generate images of the specified translation task without noticeable defects (they look realistic). Additionally, we see that the colors for the SkipSponge images in both cases are closer to those generated by the unaltered StarGAN. We further evaluate this observation through a user study as described in Section V-F.
@@
Figure 3: Examples of aging translation of the original image with various versions of the StarGAN.
@@
Figure 4: Examples of black hair translation of the original image with various versions of the StarGAN.
@@
Figure 5: Cumulative energy increases of SkipSponge after every layer on StarGAN. SkipSponge attacks for both translation tasks were done with SSIM threshold $\tau$ = 5%.
V-C Hyperparameter Study for Threshold Parameter
For SkipSponge, we perform a study on the accuracy drop thresholds with $\tau\in\{0\%,1\%,2\%,5\%\}$. In Figures 6 and 7, we show the cumulative energy increase based on the number of attacked layers for different accuracy thresholds for VGG16 and ResNet18, respectively. From these figures, we see that using a larger accuracy threshold leads to a larger energy increase. However, this is a trade-off that the attacker needs to consider, as a decrease in accuracy could make the victim suspicious of the used model. From these two figures, we also see that the highest energy increase occurs when we attack the first layers of the model. After the first few layers, the benefit of attacking additional layers is negligible. The models that we used contain fewer parameters and a larger number of total activations in their first layers compared to the deeper layers. Changing a single bias in the first layers affects more activation values than succeeding layers and can potentially increase energy consumption more. This gives SkipSponge an extra benefit, as targeting such layers makes the attack faster (fewer parameters) and more effective (more activations). We hypothesize that the negligible energy increase in the later layers is also partly due to the accuracy threshold being hit immediately after the attack has been performed on the first layer of the model, as can be seen in Figures 8 and 13.